OpenStudy (anonymous):

Find the solution set. Separate the two values with a comma: 9x^2-3x-2=0

OpenStudy (anonymous):

Using the abc-formula: \[x =3\pm \sqrt{9-4*9*-2}/18\] x=2/3 V x=-1/2